OpenStudy (anonymous):

2(x+4)+40=-4x ... can anyone solve for x

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

I am thinking it is -8

OpenStudy (anonymous):

2x + 8 + 40 = -4x 2x + 48 = -4x -2x -2x 48 = -6x /-6 /-6 x = - 8

OpenStudy (anonymous):

2(x+4)+40 = -4x 2x+8 +40 = -4x 2x+4x = -40-8 6x = -48 x= -48/6 x= -8